Scarlett is trying to decide if she should attend college or not. Part of her decision will be ba...

  1. Scarlett is trying to decide if she should attend college or not. Part of her decision will be based on the return on investment of college. She estimates that the per year cost to attend a local university in Texas including room and board is $26,346. She also assumes the costs of attendance will rise by 2.8% per year. How much is a 4-year degree going to cost Scarlett?

N

Explanation:

I/Y

PV

PMT

FV
Answer #1

First year cost = $ 26,346

Second year cost = $ 26,346 (1+2.8%) = $ 27,083.69

Second year cost = $ 26,346 (1+2.8%)^2 = $ 27,842.03

Second year cost = $ 26,346 (1+2.8%)^3 = $ 28,621.61

Total 4 year degree cost to Scarlett = $ 26,346 + $ 27,083.69 + $ 27,842.03 + $ 28,621.61 = $ 109,893.33